# A subcopula characterization of dependence for the Multivariate Bernoulli Distribution

### v1

@@ 42 Matrix{Any}:
15x7 DataFrame
 Row | variable        mean        min    median   max    nmissing  eltype
     | Symbol          Float64     Int64  Float64  Int64  Int64     DataType
   1 | SEXO            0.501803        0      1.0      1         0  Int64
   2 | TIPO_PACIENTE   0.196705        0      0.0      1         0  Int64
   3 | DIABETES        0.1348          0      0.0      1         0  Int64
   4 | EPOC            0.0113235       0      0.0      1         0  Int64
   5 | INMUSUPR        0.00843405      0      0.0      1         0  Int64
   6 | HIPERTENSION    0.174666        0      0.0      1         0  Int64
   7 | CARDIOVASCULAR  0.0158313       0      0.0      1         0  Int64
   8 | OBESIDAD        0.144864        0      0.0      1         0  Int64
   9 | RENAL_CRONICA   0.0152094       0      0.0      1         0  Int64
  10 | TABAQUISMO      0.0735159       0      0.0      1         0  Int64
  11 | e00             0.0502618       0      0.0      1         0  Int64
  12 | e20             0.38682         0      0.0      1         0  Int64
  13 | e40             0.434831        0      0.0      1         0  Int64
  14 | e65             0.128087        0      0.0      1         0  Int64
  15 | MUERTE          0.101051        0      0.0      1         0  Int64
2154719x15 Matrix{Int64}:
 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
 1  0  0  0  0  0  0  0  0  0  0  1  0  0  0
 0  0  1  0  0  0  0  0  0  0  0  0  1  0  0
 1  1  1  1  0  0  1  1  0  0  0  0  0  1  1
 0  1  1  0  0  1  0  1  0  0  0  0  1  0  1
 1  1  0  0  0  0  0  1  0  0  0  0  0  1  1
With the following description of the variables:
  1. 1.
    SEXO = $1$ for male, 00 for female.
  2. 2.
    TIPO_PACIENTE = $1$ for hospitalized, 00 non-hospitalized.
  3. 3.
    DIABETES = $1$ for diabetic, 00 if not.
  4. 4.
    EPOC = $1$ for chronic obstructive pulmonary disease (COPD), 00 if not.
  5. 5.
    INMUSUPR = $1$ for immunosuppression, 00 if not.
  6. 6.
    HIPERTENSION = $1$ for hypertension, 00 if not.
  7. 7.
    CARDIOVASCULAR = $1$ for cardiovascular disease, 00 if not.
  8. 8.
    OBESIDAD = $1$ for obesity, 00 if not.
  9. 9.
    RENAL_CRONICA = $1$ for chronic kidney disease (CKD), 00 if not.
  10. 10.
    TABAQUISMO = $1$ for smoker, 00 for non-smoker.
  11. 11.
    e00 = $1$ for ages between 00 and $19$, 00 if not.
  12. 12.
    e20 = $1$ for ages between $20$ and $39$, 00 if not.
  13. 13.
    e40 = $1$ for ages between $40$ and $64$, 00 if not.
  14. 14.
    e65 = $1$ for ages between $65$ and older, 00 if not.
  15. 15.
    MUERTE = $1$ for death, 00 if survived.
In this case $n=15$ and therefore there are $7672^{15}-1=32,$ 767 probabilities to estimate, $12^{15}-$ 1 parameters to estimate, and $162^{15}-$ 16 multivariate dependencies to estimate (pairwise, three-wise, up to 15-wise).
@time estim = MBerEst(data); # 5 minutes approx
If we are interested in the 20 largest values for dependence parameters that include variable 15 (death) then:
@@
    iMue = findall(x -> 15 in x, midx)
    [midx[iMue] m[iMue]][1:20, :]
@@
202 Matrix{Any}:
 [2, 15]              0.902144
 [2, 4, 15]           0.859008
 [2, 9, 15]           0.858144
 [2, 5, 15]           0.855051
 [2, 4, 9, 15]        0.831122
 [2, 7, 15]           0.831098
 [2, 5, 9, 15]        0.828303
 [2, 4, 5, 15]        0.827159
 [2, 7, 9, 15]        0.809365
 [2, 4, 5, 9, 15]     0.808968
 [2, 4, 7, 15]        0.808805
 [2, 5, 7, 15]        0.803182
 [2, 4, 7, 9, 15]     0.793962
 [2, 5, 7, 9, 15]     0.789702
 [2, 4, 5, 7, 15]     0.787861
 [2, 4, 5, 7, 9, 15]  0.778416
 [2, 4, 9, 14, 15]    0.659362
 [2, 4, 14, 15]       0.658259
 [2, 14, 15]          0.656512
 [2, 9, 14, 15]       0.65601
Therefore from the 20 highest values for multivariate dependence that always include death we are able to identify in all cases variable 2 (hospitalization), and in several cases variables 4, 9, 5, 7 and 14 (COPD, CKD, immunosuppression, cardiovascular disease, and age 65 or older). In this data the mortality rate was $10.1\%$ but conditioning on the mentioned variables it increases up to $66.3\%:$
begin
    pd = mean(data[:, 15])
@@
    pdcond = MBerCond(estim2.binprob.value, [7], [1,2,3,4,5,6],
                      [1,1,1,1,1,1]).binprob.dic[[1]]
@@
    println("P(death|hospital, COPD, immunosup, cardio, CKD,
             age 65+) = ", pdcond)
@@
P(death) = 0.10105122756145929
P(death|hosp,COPD,immunosup,cardio,CKD,65+) = 0.6632653061224489
6 Conclusions
In this paper, a novel subcopula-based characterization of dependence for the Multivariate Bernoulli Distribution (MBD) has been introduced. By leveraging Sklar’s theorem, it has been shown that subcopulas can effectively capture the dependence structure between Bernoulli random variables, extending the utility of subcopula theory to binary data. The proposed approach provides a flexible framework for studying multivariate binary data by separating the marginal distributions from the joint dependence structure.
Explicit formulas and bounds for both bivariate and multivariate Bernoulli distributions in terms of subcopulas were derived, allowing for a more granular understanding of the dependence parameters. In particular, dependence measures were obtained for bivariate and trivariate Bernoulli distributions, and proposed a general method to extend these results to higher dimensions. This work’s findings confirm that subcopulas are a powerful tool for modeling and quantifying dependence in the MBD of all orders, not just pairwise.
The proposed methodology not only advances the theoretical understanding of subcopulas but also opens up new possibilities for practical applications in fields dealing with multivariate binary data. Potential applications include genetics, epidemiology, and social sciences, where understanding the dependence structure between binary variables is essential. Moreover, the bayesian inference framework presented offers a robust tool for parameter estimation and uncertainty quantification, enhancing the practical implementation of the proposed approach.
Computational implementations of the MBD are quite scarce, so the set of functions provided in subsection 5.2 and downloadable from [4] is a contribution that may help in analyzing and simulating multivariate binary data. Future research could explore further generalizations of the subcopula approach to other types of discrete data and investigate its performance in applied contexts. Additionally, incorporating covariates or other forms of heterogeneity into the subcopula framework may provide more nuanced models, particularly for data sets with more complex dependence structures.

### v2

@@ Appendix A Julia code for the examples
    return accuracy
end
function predGivenGender(nsim, condProb)
    ival1 = findall(binmat[:, 1] .== 1)
    ival0 = findall(binmat[:, 1] .== 0)
    n1 = length(ival1)
    n0 = length(ival0)
    vsim = fill(-99999999, 10_000)
    accuracy = 0.0
    for k in 1:nsim
        vsim[ival1] = rand(Bernoulli(condProb.Pexit[2]), n1)
        vsim[ival0] = rand(Bernoulli(condProb.Pexit[3]), n0)
        accuracy += mean(vsim .== binmat[:, 3]) / nsim
    end
    return accuracy
end
function predGivenActive(nsim, condProb)
    ival1 = findall(binmat[:, 2] .== 1)
    ival0 = findall(binmat[:, 2] .== 0)
    n1 = length(ival1)
    n0 = length(ival0)
    vsim = fill(-99999999, 10_000)
    accuracy = 0.0
    for k in 1:nsim
        vsim[ival1] = rand(Bernoulli(condProb.Pexit[4]), n1)
        vsim[ival0] = rand(Bernoulli(condProb.Pexit[5]), n0)
        accuracy += mean(vsim .== binmat[:, 3]) / nsim
    end
    return accuracy
end
function predGivenGenderActive(nsim, condProb)
    X1X2 = Vector{Int}[]
    for j in 1:10_000
        push!(X1X2, binmat[j, [1,2]])
    end
    ival11 = findall(X1X2 .== [[1,1]])
    ival10 = findall(X1X2 .== [[1,0]])
    ival01 = findall(X1X2 .== [[0,1]])
    ival00 = findall(X1X2 .== [[0,0]])
    n11 = length(ival11)
    n10 = length(ival10)
    n01 = length(ival01)
    n00 = length(ival00)
    vsim = fill(-99999999, 10_000)
    accuracy = 0.0
    for k in 1:nsim
        vsim[ival11] = rand(Bernoulli(condProb.Pexit[6]), n11)
        vsim[ival10] = rand(Bernoulli(condProb.Pexit[7]), n10)
        vsim[ival01] = rand(Bernoulli(condProb.Pexit[8]), n01)
        vsim[ival00] = rand(Bernoulli(condProb.Pexit[9]), n00)
        accuracy += mean(vsim .== binmat[:, 3]) / nsim
    end
    return accuracy
end
# Accuracy for each rule
nsim = 1_000_000 # 4 minutes approx
@time predict=DataFrame(rule=["Non-conditional","Given gender",
                      "Given active", "Given gender & active"],
            accuracy = [predNonCond(nsim, condProb),
                        predGivenGender(nsim, condProb),
                        predGivenActive(nsim, condProb),
                        predGivenGenderActive(nsim, condProb)])
# Accuracy given gender and active status
# but setting trivariate dependence to zero
begin
    th1 = inference3v.dparam[5,3]
    th2 = inference3v.dparam[6,3]
    th3 = inference3v.dparam[7,3]
    th123 = th1 * th2 * th3
    mu12 = inference3v.dmeas[2,3]
    mu13 = inference3v.dmeas[3,3]
    mu23 = inference3v.dmeas[4,3]
    th12 = (min(th1,th2) - th1*th2)*mu12 + th1*th2
    th13 = (th1*th3 - max(th1+th3-1,0))*mu13 + th1*th3
    th23 = (th2*th3 - max(th2+th3-1,0))*mu23 + th2*th3
    p000 = th123
    p001 = th12 - th123
    p010 = th13 - th123
    p100 = th23 - th123
    p011 = th1 - th12 - th13 + th123
    p101 = th2 - th12 - th23 + th123
    p110 = th3 - th13 - th23 + th123
    p111 = 1 - (p000 + p001 + p010 + p011 + p100 + p101 + p110)
    pp = [p000, p001, p010, p011, p100, p101, p110, p111]
    Xzero3dep = MBerDep(pp)
end;
# dependence measures without and with zero trivariate dependence
[inference3v.dmeas[:, [1,3]] Xzero3dep.dmeas.value]
# Probabilities for exiting the bank with zero trivariate dependence
begin
condProb3 = DataFrame(gender = [missing,1,0,missing,missing,1,1,0,0],
                      active = [missing,missing,missing,1,0,1,0,1,0],
                      Pexit = zeros(9))
condProb3.Pexit[1] = 1 - Xzero3dep.dparam.value[7]
condProb3.Pexit[2] = MBerCond(Xzero3dep.binprob.value, [3], [1],
[1]).binprob.value[2]
condProb3.Pexit[3] = MBerCond(Xzero3dep.binprob.value, [3], [1],
[0]).binprob.value[2]
condProb3.Pexit[4] = MBerCond(Xzero3dep.binprob.value, [3], [2],
[1]).binprob.value[2]
condProb3.Pexit[5] = MBerCond(Xzero3dep.binprob.value, [3], [2],
[0]).binprob.value[2]
condProb3.Pexit[6] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
[1,1]).binprob.value[2]
condProb3.Pexit[7] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
[1,0]).binprob.value[2]
condProb3.Pexit[8] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
[0,1]).binprob.value[2]
condProb3.Pexit[9] = MBerCond(Xzero3dep.binprob.value, [3], [1,2],
[0,0]).binprob.value[2]
display(condProb3)
end
# Accuracy for each rule setting trivariate dependence to zero
nsim = 1_000_000 # 4 minutes approx
@time predict3=DataFrame(rule=["Non-conditional","Given gender",
                         "Given active","Given gender & active"],
                accuracy = [predNonCond(nsim, condProb3),
                        predGivenGender(nsim, condProb3),
                        predGivenActive(nsim, condProb3),
                        predGivenGenderActive(nsim, condProb3)]);
# Comparison of accuracies with and without trivariate dependence
println("Accuracy with trivariate dependence:")
predict
println("Accuracy with zero trivariate dependence:")
predict3
A.4 A real data example: COVID-19 in Mexico (6.4)
Ψ
Ψbegin # load data
ΨΨdf = CSV.read("covid2020.csv", DataFrame)
ΨΨshow(describe(df), allrows = true)
ΨΨdata = zeros(Int, size(df))
ΨΨfor c in 1:ncol(df)
ΨΨΨdata[:, c] = df[:, c]
ΨΨend
ΨΨprintln()
ΨΨdata
Ψend
Ψ
@time estim = MBerEst(data); # Fit MBD model (5 minutes approx)Ψ
# 20 largest values for dependence parameters
# that include variable 15 (death)
@@
iMue = findall(x -> 15 in x, midx)[midx[iMue] m[iMue]][1:20, :]
@@
Ψ
begin # conditional risk
    pd = mean(data[:, 15]) # general mortality rate
@@
    # conditional mortality rate
    pdcond = MBerCond(estim2.binprob.value, [7], [1,2,3,4,5,6],Ψ
Ψ       [1,1,1,1,1,1]).binprob.dic[[1]]
@@
    println("P(death|hospital, COPD, immunosup, cardio,
Ψ      CKD, age 65+) = ", pdcond)
@@
DECLARATIONS
Ethics approval and consent to participate
Not applicable. This study did not involve human participants, animals, or sensitive personal data requiring ethical approval.
Consent for publication
Not applicable. This manuscript does not contain data from any individual person.
Availability of data and materials
All the data and programming code is openly available at the following Github repository where the author of this repository is the same as the submitted contribution: https://github.com/aerdely/MultivariateBernoulli
Competing interests
No, I declare that the authors have no competing interests as defined by Springer, or other interests that might be perceived to influence the results and/or discussion reported in this paper.
Funding
Dirección General de Asuntos del Personal Académico, Universidad Nacional Autónoma de México, PAPIIT IN104425.
Authors’ contributions
As a single author A.E has worked all the details in this contribution.
Acknowledgements
The author gratefully acknowledges the reviewer whose comments helped to significantly improve this work.